OpenStudy (anonymous):

6x + -7y = 43 i need (x,y)

OpenStudy (skullpatrol):

What do you mean by "i need (x,y)"?

OpenStudy (skullpatrol):

(x,y) are the coordinates of a point.

OpenStudy (anonymous):

i dont know how to do this. i assume i need the points because it gives me pictures of graphs

OpenStudy (skullpatrol):

Ok let us start with graphing your equation: $$\Huge 6x +(-7y) = 43$$

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

First put it in the form: $$\Huge y= mx +b$$

OpenStudy (anonymous):

what do m and b represent

OpenStudy (skullpatrol):

The slope and y-intercept.

OpenStudy (anonymous):

and what is that?

OpenStudy (skullpatrol):

Have you seen the Khan Academy videos on YouTube about equations of lines?

OpenStudy (skullpatrol):

http://www.youtube.com/watch?v=5fkh01mClLU

OpenStudy (anonymous):

it doesnt give me slope intercept or anything

OpenStudy (anonymous):

Simplifying -6x + -7 = -43 Reorder the terms: -7 + -6x = -43 Solving -7 + -6x = -43 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '7' to each side of the equation. -7 + 7 + -6x = -43 + 7 Combine like terms: -7 + 7 = 0 0 + -6x = -43 + 7 -6x = -43 + 7 Combine like terms: -43 + 7 = -36 -6x = -36 Divide each side by '-6'. x = 6 Simplifying x = 6

OpenStudy (anonymous):

so, -36+7y=-43

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so i would do -43-(-36)=-7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7y=-7 y=-1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how would a graph a line like that?

OpenStudy (anonymous):

u would take -1,6

OpenStudy (anonymous):